Write each of the following as decimals:
70 + 8 / 10

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

70 + 8 / 10 = 70 + 0.8
= 70.8

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
70 + 8/10
Taking LCM as 10,
We get,
 
(70×10)+8/10
= 700 + 8/10
= 708/10
= 70.8
 
Hope it helps!
Thank you ✌️